# Dynamical Dark Energy from Lattice Quantum Gravity

### v2

@@ V Making contact with Cosmology
In the current epoch, where dark energy is $\approx 70\%$ of the energy density, the $\rho_{\Lambda_{0}}$ term dominates the vacuum energy density $\rho_{\Lambda}$. The $\nu$ in the prefactor of the $\rho_{\Lambda_{0}}$ term is a small correction that is not well constrained with the current precision on observations. As we run the expansion of the universe backwards in time, first $\rho_{m}$ and then eventually $\rho_{r}$ dominates the energy density of the universe. In that case $\rho_{\Lambda}\approx\nu(\rho_{m}+\rho_{r})$, which ensures that the vacuum energy density is a small fraction of the energy density of the universe, since it is suppressed by a factor of $\nu$. However, it is still much larger than $\rho_{\Lambda_{0}}$ at early times, showing that the vacuum energy density is predicted to be much larger in the early universe in this model. The current best constraints on Big Bang Nucleosynthesis (BBN) fix the Hubble scale in that era at the percent level [68]. Because the parameter $\nu$ is ${\cal O}(10^{-3})$, the corrections from this model are compatible with BBN constraints within the existing errors. Note that the logarithmic running of $\nu$ leads to an increase in $\nu$ of less than $20\%$ at the time of BBN as compared to today, so our order of magnitude estimate of $\nu$ in the late universe still applies in this earlier era.
@@ VI Conclusion and Outlook
This work presents a detailed analysis of a new set of EDT ensembles at larger volumes and finer lattice spacings than was previously possible to simulate. These ensembles have been generated using a new rejection free algorithm [19], which led to a speed-up of around a factor of one-hundred on our finest ensembles when compared with our older parallel rejection algorithm. We use this algorithm to generate several volumes across four lattice spacings. In order to meet the precision goals of this work, we introduce a new set of methods for obtaining the ratio of the direct to dual lattice spacing, $a_{\rm lat}/\ell$. Although this is a derived lattice quantity without a direct physical meaning, it is needed for constructing physical observables. The two methods introduced here to construct $a_{\rm lat}/\ell$ each assume that the emergent geometries are well-described by the classical de Sitter solution in the large volume limit. A comparison of the results from the two methods for computing $a_{\rm lat}/\ell$ show that they are in excellent agreement with each other at the percent level, thus providing a solid test that this lattice specific quantity is correctly determined.
We revisit our calculation of the absolute lattice spacing using the method introduced in Ref. [21] and a large new data set generated for this purpose. The renormalized value of Newton’s constant can be inferred from the finite-volume scaling of the tuned bare parameter $\kappa_{4}$ of the lattice action, as reviewed in Sec. III.2.1. The scaling of $\kappa_{4}$ with volume is related to Newton’s constant through a derivation assuming the Euclidean partition function is dominated by the de Sitter instanton. Using this approach to determine Newton’s constant, we are able to establish the Planck length in units of our lattice spacing. By computing Newton’s constant independently at each of our four lattice spacings, we are also able to measure the relative lattice spacing of our ensembles against a chosen fiducial spacing, defined to be the one at $\kappa_{2}=3.0$. The relative lattice spacings are in agreement with the expected behavior, with the lattices getting finer as $\kappa_{2}$ is increased. This is true for both the direct and dual lattice spacings. This result suggests that there is no barrier in principle to taking the lattice spacing arbitrarily small within the EDT formulation.
@@ Acknowledgements.
The authors thank Claude Bernard, Simon Catterall, and Scott Watson for valuable discussions and for comments on the manuscript. MD and JL were supported by the U.S. Department of Energy (DOE), Office of Science, Office of High Energy Physics under Award Number DE-SC0009998. JUY was supported by the U.S. Department of Energy grant DE-SC0019139, and by Fermi Research Alliance, LLC under Contract No. DE-AC02- 07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics, and by the Department of Energy through the Fermilab Theory QuantiSED program in the area of “Intersections of QIS and Theoretical Particle Physics.” MS acknowledges support by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. Computations were performed in part on the Syracuse University HTC Campus Grid and were supported by NSF award ACI-1341006. This research was enabled in part by support provided by ACENET, Calcul Québec, Compute Ontario, and the Digital Research Alliance of Canada. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.

### v3

@@ V Making contact with Cosmology
In the current epoch, where dark energy is $\approx 70\%$ of the energy density, the $\rho_{\Lambda_{0}}$ term dominates the vacuum energy density $\rho_{\Lambda}$. The $\nu$ in the prefactor of the $\rho_{\Lambda_{0}}$ term is a small correction that is not well constrained with the current precision on observations. As we run the expansion of the universe backwards in time, first $\rho_{m}$ and then eventually $\rho_{r}$ dominates the energy density of the universe. In that case $\rho_{\Lambda_{0}}/(\rho_{m}+\rho_{r})\ll\nu$, so that $\rho_{\Lambda}\approx\nu(\rho_{m}+\rho_{r})$. The smallness of the coefficient $\nu$ ensures that the vacuum energy density stays a small fraction of the energy density of the universe, since it is suppressed by a factor of $\nu$ compared to $\rho_{m}$ and $\rho_{r}$. However, the total vacuum energy density in the early universe is still much larger than the $\rho_{\Lambda_{0}}$ term, so that the vacuum energy density is predicted to be much larger in the early universe in this model. The current best constraints on Big Bang Nucleosynthesis (BBN) fix the Hubble scale in that era at the percent level [68]. Because the parameter $\nu$ is ${\cal O}(10^{-3})$, the corrections from this model are compatible with BBN constraints within the existing errors. Note that the logarithmic running of $\nu$ leads to an increase in $\nu$ of less than $20\%$ at the time of BBN as compared to today, so our order of magnitude estimate of $\nu$ in the late universe still applies in this earlier era.
@@ VI Conclusion and Outlook
This work presents a detailed analysis of a new set of EDT ensembles at larger volumes and finer lattice spacings than was previously possible to simulate. These ensembles have been generated using a new rejection free algorithm [19], which led to a speed-up of around a factor of one-hundred on our finest ensembles when compared with our older parallel rejection algorithm. We use this algorithm to generate several volumes across four lattice spacings. In order to meet the precision goals of this work, we introduce a new set of methods for obtaining the ratio of the direct to (renormalized) dual lattice spacing, $a_{\rm lat}/\ell$. Although this is a derived lattice quantity without a direct physical meaning, it is needed for constructing physical observables. The two methods introduced here to construct $a_{\rm lat}/\ell$ each assume that the emergent geometries are well-described by the classical de Sitter solution in the large volume limit. A comparison of the results from the two methods for computing $a_{\rm lat}/\ell$ show that they are in excellent agreement with each other at the percent level, thus providing a solid test that this lattice specific quantity is correctly determined.
We revisit our calculation of the absolute lattice spacing using the method introduced in Ref. [21] and a large new data set generated for this purpose. The renormalized value of Newton’s constant can be inferred from the finite-volume scaling of the tuned bare parameter $\kappa_{4}$ of the lattice action, as reviewed in Sec. III.2.1. The scaling of $\kappa_{4}$ with volume is related to Newton’s constant through a derivation assuming the Euclidean partition function is dominated by the de Sitter instanton. Using this approach to determine Newton’s constant, we are able to establish the Planck length in units of our lattice spacing. By computing Newton’s constant independently at each of our four lattice spacings, we are also able to measure the relative lattice spacing of our ensembles against a chosen fiducial spacing, defined to be the one at $\kappa_{2}=3.0$. The relative lattice spacings are in agreement with the expected behavior, with the lattices getting finer as $\kappa_{2}$ is increased. This is true for both the direct and renormalized dual lattice spacings. This result suggests that there is no barrier in principle to taking the lattice spacing arbitrarily small within the EDT formulation.
@@ Acknowledgements.
The authors thank Mark Alford, Claude Bernard, Simon Catterall, Jay Hubisz, Alex Maloney, and Scott Watson for valuable discussions and Mark Alford, Claude Bernard, Simon Catterall, and Scott Watson for comments on the manuscript. MD and JL were supported by the U.S. Department of Energy (DOE), Office of Science, Office of High Energy Physics under Award Number DE-SC0009998. JUY was supported by the U.S. Department of Energy grant DE-SC0019139, and by Fermi Research Alliance, LLC under Contract No. DE-AC02- 07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics, and by the Department of Energy through the Fermilab Theory QuantiSED program in the area of “Intersections of QIS and Theoretical Particle Physics.” MS acknowledges support by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. Computations were performed in part on the Syracuse University HTC Campus Grid and were supported by NSF award ACI-1341006. This research was enabled in part by support provided by ACENET, Calcul Québec, Compute Ontario, and the Digital Research Alliance of Canada. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.